# A First-Order Gradient Approach for the Connectivity Analysis of Markov Chains

### v2

@@ VIII Conclusion and future work
Assume $x\in\mathcal{X_{\varepsilon}}$ and let $J(u)$ denote the linearly (re)parameterized version of $J(x(u))$, where $P(x(u))=\sum_{i,j}x_{\ell}(u)e_{i}e_{j}^{\top}$ if $x(u)\geq 0$, with $e_{i}$ denoting the $i$th (column) basis vector, and $x(u)=A^{\dagger}b+Bu$, such that $u\in\mathbb{R}^{|E|-r}$ and $b$ as in (25). Expanding $J(u\pm\eta\Delta)$, for $u=B^{\dagger}(x-A^{\dagger}b)=B^{\top}(x-A^{\dagger}b)$, yields
$J(u\pm\eta\Delta)$ $=J(u)+\eta\sum_{i=1}^{|E|-r}\partial_{u_{i}}J(u)\Delta_{i}$
±$\pm\frac{\eta^{2}}{2}\sum_{i,j=1}^{|E|-r}\frac{\partial J(u)}{\partial_{u_{i}}\partial_{u_{j}}}\Delta_{i}\Delta_{j}$
$+\frac{\eta^{3}}{6}\sum_{i,j,k=1}^{|E|-r}\frac{\partial J(u)}{\partial_{u_{i}}\partial_{u_{j}}\partial_{u_{k}}}\Delta_{i}\Delta_{j}\Delta_{k}+\mathcal{O}(\eta^{4}).$
Then, basic algebra shows that
$\partial_{u_{l}}J(u)\approx$ $\ \frac{J(u+\eta\Delta)-J(u-\eta\Delta)}{2\eta\Delta_{l}}$
$=$ $\sum_{i=1}^{|E|-r}\partial_{u_{i}}J(u)\frac{\Delta_{i}}{\Delta_{l}}$
$+\frac{\eta^{2}}{6}\sum_{i,j,k=1}^{|E|-r}\Big{(}\frac{\partial J(y)}{\partial_{u_{i}}\partial_{u_{j}}\partial_{u_{k}}}-\frac{\partial J(y^{\prime})}{\partial_{u_{i}}\partial_{u_{j}}\partial_{u_{k}}}\Big{)}\frac{\Delta_{i}\Delta_{j}\Delta_{k}}{\Delta_{l}},$
for some $y,y^{\prime}\in[u-\eta\Delta,u+\eta\Delta]$. By assumption (A1) we have $\Delta\sim\{-1,1\}^{|E|-r}$ and by (A0) it holds that the 3rd order derivative is bounded (for a proof, use the fact that $\mathcal{X}$ is compact together with Weierstrass theorem). Then,
$\mathbb{E}_{\Delta}\left[\frac{J(u+\eta\Delta)-J(u-\eta\Delta)}{2\eta\Delta_{l}}\right]=\partial_{u_{l}}J(u)+\mathcal{O}(\eta^{2}).$
Now using that $x(u)=A^{\dagger}b+Bu$, and so
@@
where $\{x_{t}:t\geq 0\}$ is the trajectory fo the ODE, and $g(x)$ is the gradient of some function $J(x)$. Let $\tilde{g}(x)$ be a (biased) gradient estimator of $g(x)$, so that
@@
$\mathbb{E}_{\Delta}\left[\mathbb{E}_{\mathcal{E}}\left[\tilde{\delta}_{\mathcal{E}}(x(k))\right]\right]$ $=\mathbb{E}_{\Delta}\left[\mathbb{E}_{\mathcal{E}}\left[\tilde{\delta}_{\mathcal{E}}(x(k)|x(0),\ldots,x(k-1)\right]\right]$

### v4

@@ VIII Conclusion and future work
Assume $x\in\mathcal{X_{\varepsilon}}$ and let $J(u)$ denote the linearly (re)parameterized version of $J(x(u))$, where $P(x(u))=\sum_{\ell=(i,j)\in E}x_{\ell}(u)e_{i}e_{j}^{\top}$ if $x(u)\geq 0$, with $e_{i}$ denoting the $i$th (column) basis vector, and $x(u)=A^{\dagger}b+Bu$, such that $u\in\mathbb{R}^{|E|-r}$ and $b$ as in (24). Expanding $J(u\pm\eta\Delta)$, for $u=B^{\dagger}(x-A^{\dagger}b)=B^{\top}(x-A^{\dagger}b)$, yields
$J(u\pm\eta\Delta)$ $=J(u)\pm\eta\sum_{i=1}^{|E|-r}\partial_{u_{i}}J(u)\Delta_{i}$
$\hskip 11.0pt+\frac{\eta^{2}}{2}\sum_{i,j=1}^{|E|-r}\frac{\partial J(u)}{\partial_{u_{i}}\partial_{u_{j}}}\Delta_{i}\Delta_{j}$
$\hskip 11.0pt\pm\frac{\eta^{3}}{6}\sum_{i,j,k=1}^{|E|-r}\frac{\partial J(u)}{\partial_{u_{i}}\partial_{u_{j}}\partial_{u_{k}}}\Delta_{i}\Delta_{j}\Delta_{k}+\mathcal{O}(\eta^{4}).$
We can use the expansion to estimate partial derivatives $\partial_{u_{l}}J(u)$, for any $l$, as we show in the following. Basic algebra shows that for a given $\Delta\in\{-1,1\}^{d}$, and some $y,y^{\prime}\in[u-\eta\Delta,u+\eta\Delta]$;
$\begin{split}&\hskip-15.0pt\frac{J(u+\eta\Delta)-J(u-\eta\Delta)}{2\eta\Delta_{l}}=\\ &=\sum_{i=1}^{|E|-r}\partial_{u_{i}}J(u)\frac{\Delta_{i}}{\Delta_{l}}\\ &\hskip 11.0pt+\frac{\eta^{2}}{6}\sum_{i,j,k=1}^{|E|-r}\hskip-2.0pt\big{(}\frac{\partial J(y)}{\partial_{u_{i}}\partial_{u_{j}}\partial_{u_{k}}}-\frac{\partial J(y^{\prime})}{\partial_{u_{i}}\partial_{u_{j}}\partial_{u_{k}}}\big{)}\hskip-1.0pt\frac{\Delta_{i}\Delta_{j}\Delta_{k}}{\Delta_{l}}.\end{split}$
Note that by (A0) it holds that the 3rd order derivative is bounded (for a proof, use the fact that $\mathcal{X}$ is compact together with Weierstrass theorem). Then, taking the expectation over $\Delta\sim\text{Rad}^{d}$ (i.e., each $\Delta_{i}$ i.i.d. uniform over $\{-1,1\}$) by assumption (A1), we obtain:
$\mathbb{E}_{\Delta}\left[\frac{J(u+\eta\Delta)-J(u-\eta\Delta)}{2\eta\Delta_{l}}\right]=\partial_{u_{l}}J(u)+\mathcal{O}(\eta^{2}),$
i.e., effectively estimating $\partial_{u_{l}}J(u)$. Now using that $x(u)=A^{\dagger}b+Bu$, and so
@@
where $\{x_{t}:t\geq 0\}$ is the trajectory of the ODE, and $g(x)$ is the gradient of some function $J(x)$. Let $\tilde{g}(x)$ be a (biased) gradient estimator of $g(x)$, so that
@@
$\mathbb{E}_{\Delta}\hskip-1.0pt\left[\mathbb{E}_{\mathcal{E}}\hskip-1.0pt\left[\tilde{\delta}_{\mathcal{E}}(x(k))\right]\right]\hskip-1.0pt$ $=\mathbb{E}_{\Delta}\hskip-1.0pt\left[\mathbb{E}_{\mathcal{E}}\hskip-1.0pt\left[\tilde{\delta}_{\mathcal{E}}(x(k)|x(0),\ldots,x(k-1)\right]\right]$